OpenStudy (anonymous):

All 37 students in a math class attended on Monday. On Tuesday 25 students attended class. What was the percent of decrease? Round to two decimal points. show work

OpenStudy (anonymous):

it says the answer is A

OpenStudy (anonymous):

32.40%

OpenStudy (anonymous):

its 37-25=12/37=.3243*100= 32.40%

OpenStudy (anonymous):

My bad, \[ \frac {37-25}{37}\times 100 \% =\frac{1200}{37}\% =32.4324 \% \]